# Genetic Instruct: Scaling up Synthetic Generation of Coding Instructions for Large Language Models

### v2

@@ Abstract
Somshubra Majumdar*, Vahid Noroozi thanks: Equal contribution, Mehrzad Samadi, Sean Narenthiran, Aleksander Ficek, Wasi Uddin Ahmad, Jocelyn Huang, Jagadeesh Balam, Boris Ginsburg NVIDIA Santa Clara, CA 15213, USA {smajumdar,vnoroozi,msamadi,snarenthiran,aficek, wasiuddina,jocelynh,jbalam,bginsburg}@nvidia.com
@@ 1 Introduction
Large Language Models (LLMs) have made significant progress in programming tasks and are increasingly used as code assistants (Liang et al., 2024). To fully exploit their potential, they require alignment (Ouyang et al., 2022), which depends on paired instruction-solution examples to shape the behavior of the model. However, creating diverse and complex instructions, especially in coding domains, can be expensive due to the need for expert input. A promising alternative is generating synthetic instructions using another LLM. Prior research shows that synthetic instructions are effective for both coding (Luo et al., 2024; Wu et al., 2024; Wei et al., 2024b; Yu et al., 2024) and general tasks (Wang et al., 2023; Honovich et al., 2023; Xu et al., 2024).
@@
In this paper, we introduce Genetic-Instruct, a scalable algorithm for generating synthetic coding instructions, illustrated in Figure 1. Inspired by evolutionary algorithms, Genetic-Instruct starts with a small set of seed instructions and uses LLMs to generate new instruction-code pairs through two operations of crossover and mutation.
The crossover operation follows a self-instruct approach Wang et al. (2023), where an LLM creates new instructions from few-shot examples, expanding topic coverage beyond the original seeds. The crossover operator is mainly employed to enhance diversity by expanding the overall coverage of the instructions to wider domains and topics beyond the original seed instructions.
In the mutation operation, an LLM evolves a given instruction into another instruction based on some predefined rules (Luo et al., 2024). This operation can help the generation process to increase the diversity of the instructions locally. An instruction generated by one operation is added to the pool of the seeds and it may be be used by the the operation or other in the next step. This collaborative and coupled interaction between the crossover and mutation is the main key foundation of our proposed approach. It boosts instruction diversity which is an essential factor in the success of synthetic instruction generation.
@@
Additionally, the entire pipeline is designed for efficient parallel execution with multiple colonies of populations by running multiple instances of this process in parallel. Furthermore, this process can get repeated multiple times to generate more generations by using the instructions generated from previous round as the seed for the next generation.
Using our Genetic-Instruct algorithm, we generated a large dataset of synthetic coding instructions (more than 7.5M samples), starting from 512 seed questions. We trained LLMs on this data via supervised fine-tuning (SFT) and evaluated them on code generation benchmarks. Our work supports open-source development, avoiding any closed-source data or models.
Models trained on our synthetic dataset achieved strong results across coding benchmarks, outperforming other instruction generation methods and also some of the existing public SFT datasets. Our experiments also show that Genetic-Instruct can produce high-quality data without requiring very strong LLMs or large seed sets.
@@ 3 Genetic-Instruct
Subsequently, a decontamination process is applied to minimize risk of contaminated instructions in the training data. The complete pipeline is illustrated in Figure 1 for one generation, and the procedure for the whole algorithm is detailed in Algorithm 1. In the following section, each step is explained in detail.
@@ 3.6 LLM Decontamination
To prevent any evaluation benchmark questions from leaking into our training samples, we adopted the decontamination methodology proposed by Yang et al. (2023), which involves two primary stages. First, for each synthesized question, we performed an embedding-based similarity search using a Sentence Transformer Reimers and Gurevych (2020) model to identify the most similar test example from all benchmark datasets. Second, we constructed question pairs by matching each synthesized question with its most similar test example. An LLM, specifically Meta-Llama-3-70B-Instruct, was then employed to evaluate whether any of these pairs constituted a paraphrase (details on the prompt are provided in Appendix  7). To control for potential positional bias in the LLM’s paraphrase detection, we generated two pairs for each match: one where the synthesized question appeared first and another where the test set question was presented first (Toshniwal et al., 2024). If any of these pairs were determined to be similar by the LLM, the synthesized question was removed.
@@ 4.1 Experimental Settings
Gen. Algo. / Dataset $|S|$ MBPP MBPP+ HE HE+ Avg
@@
Table 1: Comparison of Genetic-Instruct with other data generation algorithms and datasets. $|S|$ indicates numbers of training samples.
@@ 4.2 Performance Evaluation
In this section, we evaluate the effectiveness of our proposed approach for generating synthetic supervised fine-tuning (SFT) samples aimed at enhancing the coding capabilities of large language models (LLMs). We used Llama3.1-8B-Base (Grattafiori et al., 2024) as the base model and employed Mixtral-8x22B (Jiang et al., 2024) as the Instructor-LLM, Coder-LLM, and Judge-LLM.
@@
4.3 Ablation Study
In this ablation study, we assess the impact of mutation and crossover operations in Genetic-Instruct on the quality of generated data. We compare three setups: Crossover-Only, where only the crossover operation is used during data generation; Mutation-Only, where only the mutation operation is applied; and the full Genetic-Instruct approach, which employs both. For each, we generated three generations totaling 4 million samples and fine-tuned a Llama3.1-8B Base model to evaluate downstream performance. This setup allows us to assess the individual and combined impact of these genetic operators on downstream model performance. Mutation-Only resembles WizardCoder conceptually, but with a key distinction: it updates the evolving seed pool with newly generated samples, unlike WizardCoder, which evolves only the initial seeds.
@@ 4.3 Ablation Study
4.4 Influence of the Generator Model
Base Model Generation Model MBPP MBPP+ HE HE+ Avg
@@ 4.4 Influence of the Generator Model
Table 3: Ablation study on the effect of the generator model on data generation quality.
@@ 5 Conclusion
We introduced Genetic-Instruct, a novel algorithm inspired by evolutionary principles for generating synthetic coding instructions for large language models. Genetic-Instruct is specifically designed to support parallel generation, making it a highly scalable solution for synthetic data creation. We benchmarked our approach against several baseline methods and publicly available datasets, and the results consistently demonstrated its effectiveness in improving performance on downstream code generation tasks.

### v3

@@ Abstract
Somshubra Majumdar*, Vahid Noroozi thanks: Equal contribution, Mehrzad Samadi, Sean Narenthiran, Aleksander Ficek, Wasi Uddin Ahmad, Jocelyn Huang, Jagadeesh Balam, Boris Ginsburg NVIDIA {smajumdar,vnoroozi,msamadi,snarenthiran,aficek, wasiuddina,jocelynh,jbalam,bginsburg}@nvidia.com
@@ 1 Introduction
Large Language Models (LLMs) have made significant progress in programming tasks and are increasingly being used as code assistants (Liang et al., 2024). To fully exploit their potential, they require alignment (Ouyang et al., 2022), which depends on paired instruction-solution examples to shape the behavior of the model. However, creating diverse and complex instructions, especially in coding domains, can be expensive due to the need for expert input. A promising alternative is to generate synthetic instructions using another LLM. Previous research shows that synthetic instructions are effective for both coding (Luo et al., 2024; Wu et al., 2024; Wei et al., 2024b; Yu et al., 2024) and general tasks (Wang et al., 2023; Honovich et al., 2023; Xu et al., 2024).
@@
In this paper, we introduce Genetic-Instruct, a scalable algorithm to generate synthetic coding instructions, illustrated in Figure 1. Inspired by evolutionary algorithms, Genetic-Instruct starts with a small set of seed instructions and uses LLMs to generate new instruction-code pairs through two operations of crossover and mutation.
The crossover operation follows a self-instruct approach Wang et al. (2023), where an LLM creates new instructions from few-shot examples, expanding the topic coverage beyond the original seeds. The crossover operator is mainly employed to enhance diversity by expanding the overall coverage of the instructions to wider domains and topics beyond the original seed instructions.
In the mutation operation, an LLM evolves a given instruction into another instruction based on some predefined rules (Luo et al., 2024). This operation can help the generation process to increase the diversity of the instructions locally. An instruction generated by one operation is added to the pool of the seeds, and it may be used by the the operation or other in the next step. This collaborative and coupled interaction between the crossover and mutation is the main key foundation of our proposed approach. It boosts instruction diversity, which is an essential factor in the success of synthetic instruction generation.
@@
Additionally, the entire pipeline is designed for efficient parallel execution with multiple colonies of populations by running multiple instances of this process in parallel. Furthermore, this process can be repeated multiple times to generate more generations using the instructions generated from the previous round as the seed for the next generation.
Using our Genetic-Instruct algorithm, we generated a large dataset of synthetic coding instructions (more than 7.5M samples), starting from 512 seed questions. We trained LLMs on these data via supervised fine-tuning (SFT) and evaluated them on code generation benchmarks. Our work supports open-source development, avoiding any closed-source data or models.
Models trained on our synthetic dataset achieved strong results across coding benchmarks, outperforming other instruction generation methods and also some of the existing public SFT datasets. Our experiments also show that Genetic-Instruct can produce high-quality data without requiring very strong LLMs or large seed sets. We released the dataset publicly to support open-source LLM development  [^1] https://huggingface.co/datasets/nvidia/OpenCodeGeneticInstruct.
@@ 3 Genetic-Instruct
Subsequently, a decontamination process is applied to minimize risk of contaminated instructions in the training data. The complete pipeline is illustrated in Figure 1 for one generation, and the procedure for the whole algorithm is detailed in Algorithm 1. In the following, each step is explained in detail.
@@ 3.6 LLM Decontamination
To prevent any evaluation benchmark questions from leaking into our training samples, we adopted the decontamination methodology proposed by Yang et al. (2023), which involves two primary stages. First, for each synthesized question, we performed an embedding-based similarity search using a Sentence Transformer Reimers and Gurevych (2020) model to identify the most similar test example from all benchmark datasets. Second, we constructed question pairs by matching each synthesized question with its most similar test example. An LLM, specifically Meta-Llama-3-70B-Instruct, was then employed to evaluate whether any of these pairs constituted a paraphrase (details on the prompt are provided in Appendix  7).
To control for potential positional bias in the LLM’s paraphrase detection, we generated two pairs for each match: one where the synthesized question appeared first and another where the test set question was presented first (Toshniwal et al., 2024). If any of these pairs were determined to be similar by the LLM, the synthesized question was removed.
@@ 4.1 Experimental Settings
Generation Algorithm/Dataset Data Size MBPP MBPP+ HumanEval HumanEval+ Average
@@
Table 1: Comparison of Genetic-Instruct with other data generation algorithms and datasets. Average of the accuracies on all the benchmarks are also reported.
@@ 4.2 Performance Evaluation
In this section, we evaluate the effectiveness of our proposed approach for generating synthetic supervised fine-tuning (SFT) samples aimed at enhancing the coding capabilities of LLMs. We used Llama3.1-8B-Base (Grattafiori et al., 2024) as the base model and employed Mixtral-8x22B (Jiang et al., 2024) as the Instructor-LLM, Coder-LLM, and Judge-LLM.
@@
Base Model Generation Model MBPP MBPP+ HumanEval HumanEval+ Average
@@
Table 3: Ablation study on the effect of the generator model on the quality of the data generation. Average of the accuracies on all the benchmarks are also reported.
4.3 Ablation Study
In this ablation study, we assess the impact of mutation and crossover operations in Genetic-Instruct on the quality of generated data. We compare three setups: Crossover-Only, where only the crossover operation is used during data generation; Mutation-Only, where only the mutation operation is applied; and the full Genetic-Instruct approach, which employs both.
For each setup, we generated three generations totaling 4 million samples and fine-tuned a Llama3.1-8B Base model to evaluate downstream performance. This setup allows us to assess the individual and combined impact of these genetic operators on downstream model performance. Mutation-Only resembles WizardCoder conceptually, but with a key distinction: it updates the evolving seed pool with newly generated samples, unlike WizardCoder, which evolves only the initial seeds.
4.4 Influence of the Generator Model
@@ 5 Conclusion
We introduced Genetic-Instruct, a novel algorithm inspired by evolutionary principles to generate synthetic coding instructions for LLMs. Genetic-Instruct is specifically designed to support parallel generation, making it a scalable solution for synthetic data creation. We benchmarked our approach against several baseline methods and publicly available datasets, and the results consistently demonstrated its effectiveness in improving performance on code generation tasks. Also in our ablation studies, we demonstrated the effectiveness of combining the two main operations to achieve the best performance. We publicly released the 7.5M synthetic instruction-solution dataset to facilitate the development of open source LLMs.